OpenStudy (anonymous):

Factoring cubes/Solve by factoring: How do I do this when it is a problem that is to the 4th power? X^4-625

OpenStudy (anonymous):

What number times 4 will give -625? You could go through polynomial division but its not worth it because its a perfect quartic.

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Exactly, so how does that look once you reduce both parts by the fourth power?

OpenStudy (anonymous):

x-5

OpenStudy (anonymous):

Ill be honest with you and say that I can't 100% continue on from here. Could someone else take over?

zepdrix (zepdrix):

rosey, we'll have to do something similar that we did in #2 of the last problem set. Remember how we messed with the power? We don't have a way of reducing 4th powers, but do you remember a formula for the difference of SQUARES? :)

OpenStudy (anonymous):

It's ok. Thanks for trying

OpenStudy (anonymous):

No I do not refresh my memory please

OpenStudy (tkhunny):

"What number times 4 will give -625" ?? -625 / 4 = 156.25 Try this: What number raised to the fourth power is 625? 5^4 = 625 Somehow, you understood that and got it right. Kind of amazing.

zepdrix (zepdrix):

\[\large a^2-b^2=(a-b)(a+b)\] We want to utilize this formula :)

OpenStudy (anonymous):

Ok.

zepdrix (zepdrix):

\[\huge x^4=x^{2*2}\] Remember how we rewrote that as a lower power last time? :D

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

I was trying to do that on my own but it wouldn't add up to the 4th power when putting into the cube formula. I guess I was using the wrong formula?

zepdrix (zepdrix):

you could apply the cube formula, but you would have fractional exponents, it would look really really ugly! :) Because you would rewrite:\[\large x^4=x^{(4/3)3}=(x^{4/3})^3\]

OpenStudy (anonymous):

oh no - I don't want that. It looks ugly

zepdrix (zepdrix):

We have really nice formulas for the difference of SQUARES, and difference of CUBES. When you look at your problem, try to identify if the power is a multiple of 2 or 3, that might help you figure out which rule to use.

OpenStudy (anonymous):

multiple of 2

OpenStudy (anonymous):

I got it

zepdrix (zepdrix):

\[\large x^4-625=x^4-5^4=x^{2*2}-5^{2*2}=(x^2)^2-(5^2)^2\] Ok understand what's going on now? :D yay

OpenStudy (anonymous):

\[(x ^{2}+25)(x+5)(x-5)\]

zepdrix (zepdrix):

Ok good good :) So you recognized the difference of squares in your answer, so you broke it down another time. yay rosey! \c:/

OpenStudy (anonymous):

YEAYYYY. Cool thank you so much!!!!